OpenStudy (anonymous):

It takes Mary 45 minutes to drive 30 miles. It takes John 3.5 hours to drive 200 miles. Who is driving faster? How much faster are they driving? *only help set up the equation not the answer*

OpenStudy (aum):

45 minutes is 3/4 hour. Mary's speed = 30 / (3/4) mph = ? John's speed = 200 / 3.5 mph = ?

OpenStudy (anonymous):

This is essentially like your last question. First find out how many miles per minute Mary is driving, then do the same with John, but this time with John you must convert hours to minutes first